# EMMA: Efficient Visual Alignment in Multi-Modal LLMs

### v1

@@ Abstract
Multi-modal Large Language Models (MLLMs) have recently exhibited impressive general-purpose capabilities by leveraging vision foundation models to encode the core concepts of images into representations. These are then combined with instructions and processed by the language model to generate high-quality responses. Despite significant progress in enhancing the language component, challenges persist in optimally fusing visual encodings within the language model for task-specific adaptability. Recent research has focused on improving this fusion through modality adaptation modules but at the cost of significantly increased model complexity and training data needs. In this paper, we propose EMMA (Efficient Multi-Modal Adaptation), a lightweight cross-modality module designed to efficiently fuse visual and textual encodings, generating instruction-aware visual representations for the language model. Our key contributions include: (1) an efficient early fusion mechanism that integrates vision and language representations with minimal added parameters (less than 0.2% increase in model size), (2) an in-depth interpretability analysis that sheds light on the internal mechanisms of the proposed method; (3) comprehensive experiments that demonstrate notable improvements on both specialized and general benchmarks for MLLMs. Empirical results show that EMMA boosts performance across multiple tasks by up to 9.3% while significantly improving robustness against hallucinations. Our code is available at https://github.com/SaraGhazanfari/EMMA
@@ 1 Introduction
Figure 1: This figure presents the EMMA’s Architecture where Modality Adaptation is introduced to the standard MLLM architecture. The Modality Adaptation module consists of the Text Encoder, Instruction Projection, and Visual alignment modules. EMMA enhances the multi-modality alignment by incorporating the instruction encodings generated by the Text Encoder which are then projected to the joint space by the Instruction Projection module. The concatenating visual and textual encodings are then passed through the Visual Alignment module, resulting in multi-modal, instruction-aware representations.
@@
Over the past years, Large Language Models (LLMs) have transformed natural language processing (NLP) by demonstrating exceptional abilities in understanding, generating, and reasoning with text across a wide range of tasks; from machine translation and summarization to complex problem-solving and conversational agents (Touvron et al., 2023; Zheng et al., 2023). However, many real-world applications require the ability to process more than just text, such as understanding visual content or synthesizing information from different modalities (Ghazanfari et al., 2023a; b). This has led to the development of multi-modal LLMs  [^1] In this paper, we refer use multi-modality specifically for combination of image and text modalities., which combine the linguistic strengths of LLMs with vision foundation models, enabling cross-modal understanding and reasoning. By integrating textual and visual information, these models extend the capabilities of traditional LLMs to address tasks like image captioning, visual question answering, and text-to-image generation(Liu et al., 2024b; Alayrac et al., 2022; Achiam et al., 2023).
@@
Current state-of-the-art multi-modal models typically rely on fixed visual feature encodings extracted from vision foundation models, which are projected into the text space and passed to the language model along with instructions (Driess et al., 2023b; Liu et al., 2024a; b). However, the static nature of these encodings, formed without considering the instruction, limits the model’s ability to adapt dynamically to specific tasks or contexts. This disconnect between visual and textual components reduces flexibility, making the model less responsive to task-specific nuances. To address this, BLIP-2 (Li et al., 2023b) introduced a cross-attention-based module (called Q-former) to integrate the visual and instruction encodings, a design later adopted by others (Li et al., 2023b; Huang et al., 2023; Ye et al., 2023b). The current state-of-art model, mPLUG-Owl2 (Ye et al., 2024), introduces a modality adaptation module that employs attention modules to embed the two modalities into a shared semantic space and thus enhances cross-modality collaboration. mPLUG-Owl2’s performance improvement comes with several limitations compared to baseline models. First, mPLUG-Owl2 leverages LLaMA’s text embeddings and CLIP’s vision encoder to generate the instruction and visual encodings respectively. Therefore the encodings are generated using two distinct models with no initial multi-modality alignment. Second, mPLUG-Owl2’s modality-adaptive module introduces roughly 1B more parameters, 3×3$\times$3 × more than its vision encoder. The modularity adaptation module is then trained from scratch, requiring 348 million image-text pairs for pertaining, 300×300$\times$300 × more than the baseline. Third, the vision encoder requires training during both the pretraining and instruction-tuning stages, which increases the overall training cost and makes the vision encoder more susceptible to loss of generality. Finally, except for a few benchmarks, the model offers only marginal improvements, and in some cases, performs worse than the baseline.
@@
  • EMMA Outperforms Larger Models: Compared to mPLUG-Owl2 which has 50×50$\times$50 × larger modality adaptation module and is trained on 300×300$\times$300 × more data, EMMA outperforms on 7 of 8 benchmarks. Additionally, compared with BRAVE which has 24×24$\times$24 × larger vision encoder and is trained on 100×100$\times$100 × more data EMMA outperforms on all benchmarks.
@@ 2 Related Work
In recent years, there has been significant progress in the development of multi-modal large language models (LLMs) that integrate vision and language to handle tasks requiring both modalities (Zhang et al., 2024; 2023; Wu et al., 2023; Sun et al., 2024; Alayrac et al., 2022; Lai et al., 2023; Li et al., 2023a; g; d; Lin et al., 2024; Liu et al., 2023b; c; Tian et al., 2024; Wang et al., 2024b; c; Chen et al., 2023a). By combining the language understanding of LLMs with the perceptual abilities of vision foundation models, multi-modal LLMs are able to tackle a wide array of tasks that require cross-modal alignment and understanding. We can divide these models into two general categories based on the way that the two visual and textual modalities are integrated. In the first category, including LLaVA (Liu et al., 2024b; a), PaLM-E (Driess et al., 2023a), Shikra (Chen et al., 2023b), etc., the vision encodings are projected to the text space with a few linear layers, and then concatenated with the instruction tokens and passed to the LLMs. In the second category, a more complex module is used for cross-modality adaptation, where both visual and textual encodings are processed through the adaptation module. First introduced with Flamingo (Alayrac et al., 2022) and later adopted by BLIP-2 (Li et al., 2023c), InstructBLIP (Dai et al., 2024), Qwen-VL (Bai et al., 2023), mPLUG-Owl Ye et al. (2023a), and MiniGPT-4 (Zhu et al., 2024), the use of a Q-former (a cross-attention based module) has become a prominent technique.
@@
InstructBLIP 1B ×$\times\text{3.3}$ 0.2B×$\times\text{10}$ 7B 7.91B
BLIP-2/InstructBLIP 1B×$\times\text{3.3}$ 0.2B×$\times\text{10}$ 13B 14.2B
Qwen-VL-Chat 1.9B×$\times\text{6.3}$ 0.08B×$\times\text{4}$ 7.7B 9.6B
BRAVE 7.2B×$\times\textbf{24}$ 0.1B×$\times\text{5}$ 2.6B 10B
mPLUG-Owl2 0.3B×$\times\text{1}$ 1B×$\times\textbf{50}$ 7B 8.2B
EMMA 0.3B 0.02B 7B 7.3B
Table 1: Paremeter sizes of EMMA compared with state-of-art MLLMs across three components: Vision Encoder, Modality Adapter, and LLM. EMMA uses the same-sized Vision Encoder and LLM as mPLUG-Owl2 but its Modality Adapter, our main contribution, is 50$\times$ smaller.
@@
Since the emergence of multi-modal LLMs, aligning visual and textual modalities has remained a significant challenge in achieving robust and seamless integration. Previous works (Alayrac et al., 2022; Li et al., 2023c; Ye et al., 2023a; c; Kar et al., 2024) have predominantly focused on utilizing Q-formers and similar cross-attention modules as modality adaptation components to integrate visual and textual embeddings. Recent advancements in modality adaptation include mPLUG-Owl2 (Ye et al., 2023c) and BRAVE (Kar et al., 2024). mPLUG-Owl2 (Ye et al., 2023c) introduces 1B modality adaptation module that employs distinct parameters to project multiple modalities into a unified semantic space for enhanced modality adaptation. On the other hand, BRAVE (Kar et al., 2024) leverages a concatenation of various visual encodings, summing up to 7B, directly feeding into the Q-former. These modality adaptation modules rely on intricate architectures, introducing millions to billions of additional parameters to the model, which significantly increases the computational costs. This added complexity not only demands vast amounts of training data but also imposes considerable overhead during inference. Furthermore, the complication of these systems makes it difficult to discern the primary drivers behind performance improvements — whether they stem from the model complexity, the early fusion of vision and text encodings, or the sheer volume of additional training data. This leads us to the central focus of our work: (1) addressing the inefficiencies in current multi-modal models by proposing a more streamlined approach. Our goal is to achieve an efficient early fusion of visual and textual encodings without significantly increasing the number of parameters or computational costs. (2) Conducting a thorough analysis to determine the key drivers behind performance improvements and the dynamics of the modality adaptation module to generate the multi-modal encodings. This investigation will help isolate the most effective factors and guide future advancements in multi-modal LLMs.
@@
The evaluation of multi-modal LLMs relies on a mix of traditional academic benchmarks and newer ones tailored to instruction-following MLLMs. Established benchmarks like VQA-v2 (Goyal et al., 2017) and GQA Hudson & Manning (2019) gauge a model’s ability to interpret visuals through open-ended, short-answer questions. ScienceQA (Lu et al., 2022) tests zero-shot generalization in scientific question answering, while VizWiz (Gurari et al., 2018) offers real-world images captured by visually impaired users, challenging models with poor framing, blur, and other quality issues typical of non-professional photos. Additionally, newer benchmarks target instruction-following MLLMs. MathVista (Lu et al., 2023) introduces diverse challenges from mathematical and visual tasks. MMMU (Yue et al., 2024) evaluates multi-modal models on a broad range of college-level tasks requiring deep subject knowledge and reasoning. To assess multi-image reasoning, MUIRBENCH (Wang et al., 2024a) provides a comprehensive benchmark with 12 diverse tasks to evaluate the multi-image understanding abilities of MLLMs.
For general robustness, MMBench (Liu et al., 2023a) offers a multiple-choice visual question answering benchmark in both English and Chinese, suggesting shuffling the choices to test the model’s robustness to the order of options. MMVP (Tong et al., 2024) evaluates robustness by identifying similar images with minute differences and manually pinpointing the visual details the CLIP vision encoder overlooks, which leads to incorrect responses from MLLMs. For hallucination, POPE (Li et al., 2023f) examines the extent of hallucinations across three COCO subsets. AMBER (Wang et al., 2023), a multi-dimensional benchmark, evaluates generative and discriminative tasks, including existence, attribute, and relation hallucination. FOIL (Shekhar et al., 2017) pairs images with both accurate and ”foil” captions—captions that are nearly identical but contain a single error—further challenging the model’s accuracy.
(a) Histogram of $\ell_{2}$ Norms
(b) Histogram of $\ell_{1}$ Norms
Figure 2: To evaluate the contribution of visual and textual tokens to the Aligned Representations, we analyze the weight matrix of the Visual Alignment module, which consists of a linear layer. The histogram displaying the $\ell_{1}$ and $\ell_{2}$ norms of the weights for each token is presented in this figure, revealing that visual tokens have a stronger influence on the aligned representations. Additionally, the impact of textual tokens varies, prompting further investigation in Figure 5(a) to identify which specific textual tokens contribute more significantly to the final representations.
@@ 3.1 EMMA: Efficient Multi-Modal Adaptation
Figure 3: Comparing EMMA with the state-of-the-art model LLaVA-1.5 demonstrates that EMMA surpasses the baseline across a range of benchmarks up to 9.3%.
Recent progress in multi-modal models is largely driven by the robust reasoning capabilities of large language models. Therefore, a persistent challenge remains in effectively aligning these two modalities to ensure seamless fusion and task-specific adaptability. Current approaches often depend on complex cross-modality modules, which introduce a significant number of parameters and thus require large amounts of training data. We hypothesize that the need for a complex modality adaptation module arises from the fact that visual and textual encodings are produced by two independently trained independently that are themselves unaligned. This is precisely the case for mPLUG-Owl2, which utilizes CLIP as its vision encoder and LLaMA’s text embeddings for its text encoder. As a result, the multi-modality module must, in addition to incorporating text information into the visual embedding, also first align the two embeddings.
@@
We refer to our proposed architecture as EMMAEfficient Multi-Modal Adaptation. Figure 1 illustrates EMMA’s architecture. The left side highlights the high-level structure, where the standard modules of a multi-modal LLM are shaded in gray, while EMMA’s newly introduced Modality Adaptation module is shown in green. On the right, the details of the Modality Adaptation module are depicted. CLIP’s text encoder is used to generate instruction encodings, which are then processed through the Instruction Projection module to prepare them for fusion with the visual tokens. Once the visual and textual representations are generated, we introduce early fusion via a lightweight module called the Visual Alignment module. This component, consisting of a simple linear layer, combines visual and textual tokens to create the model’s multi-modal encoding. The dimensions of this alignment layer are designed to maintain the same number of visual tokens as the baseline model, ensuring consistency in the number of visual tokens passed to the language model. The Visual Alignment module plays a critical role in ensuring effective alignment between visual and instructions encodings, highlighting the most relevant tokens from the visual encodings in response to the instruction, and thereby delivering more precise visual information to the language model. Its lightweight design facilitates easier interpretability and analysis and more importantly, reduces the training and inference time while outperforming the state-of-the-art models with 10×10$\times$10 × less parameters. Additionally, their initial alignment allows the text and vision encoders to remain frozen during training, preventing the model from overfitting to the training data.
For training, we follow the same two-stage instruction fine-tuning process as LLaVA. In the pretraining stage, only the Visual Alignment and Projection modules are trained, while the language model remains frozen. During the fine-tuning stage, the LLM is unfrozen and fine-tuned along with the two aforementioned modules. To ensure that any performance improvements are not simply due to the addition of more data, we use the same dataset as the baseline model. Figure 3 demonstrates the improvements achieved by our proposed method, EMMA, across several benchmarks compared to the baseline model, LLaVA-1.5. EMMA consistently outperforms LLaVA-1.5 on tasks such as MMVP (Tong et al., 2024) (+9.3%), MuirBench (Wang et al., 2024a)(+5.62%), SQA (Lu et al., 2022) (+4.2%), MMBench (Liu et al., 2023a) (+1.9%), and POPE (Li et al., 2023e) (+1.17%) showing consistent gains in cross-modal tasks requiring visual and textual understanding.
(a) $\ell_{1}$ norm of weights associated with textual tokens.
(b) $\ell_{2}$ distance between the MMVP image pairs.
Figure 5: Left: $\ell_{1}$ Norm of Weights Associated with Textual Tokens. The bar plot reveals that the early tokens are assigned higher weights in the Visual Alignment module, placing greater emphasis on them. This aligns with the typical configuration of MLLMs, where instructions are frequently given as concise prompts. Right: $\ell_{2}$ distance between the MMVP image pairs. This figure demonstrates the reason behind EMMA’s relative success in differentiating between similar images compared with LLaVA.

### v2

@@ Abstract
Multi-modal Large Language Models (MLLMs) have recently exhibited impressive general-purpose capabilities by leveraging vision foundation models to encode the core concepts of images into representations. These are then combined with instructions and processed by the language model to generate high-quality responses. Despite significant progress in enhancing the language component, challenges persist in optimally fusing visual encodings within the language model for task-specific adaptability. Recent research has focused on improving this fusion through modality adaptation modules, but at the cost of significantly increased model complexity and training data needs. In this paper, we propose EMMA (Efficient Multi-Modal Adaptation), a lightweight cross-modality module designed to efficiently fuse visual and textual encodings, generating instruction-aware visual representations for the language model. Our key contributions include: (1) an efficient early fusion mechanism that integrates vision and language representations with minimal added parameters (less than 0.2% increase in model size), (2) an in-depth interpretability analysis that sheds light on the internal mechanisms of the proposed method; (3) comprehensive experiments that demonstrate notable improvements on both specialized and general benchmarks for MLLMs. Empirical results show that EMMA boosts performance across multiple tasks by up to 9.3% while significantly improving robustness against hallucinations.
@@ 1 Introduction
Over the past years, Large Language Models (LLMs) have transformed natural language processing (NLP) by demonstrating exceptional abilities in understanding, generating, and reasoning with text across a wide range of tasks; from machine translation and summarization to complex problem-solving and conversational agents (Touvron et al., 2023; Zheng et al., 2023). However, many real-world applications require the ability to process more than just text, such as understanding visual content or synthesizing information from different modalities Ghazanfari et al. (2023b; a). This has led to the development of multi-modal LLMs  [^1] In this paper, we refer use multi-modality specifically for the combination of image and text modalities., which combine the linguistic strengths of LLMs with vision foundation models, enabling cross-modal understanding and reasoning. By integrating textual and visual information, these models extend the capabilities of traditional LLMs to address tasks like image captioning, visual question answering, and text-to-image generation(Liu et al., 2024b; Alayrac et al., 2022; Achiam et al., 2023).
@@
Current state-of-the-art multi-modal models typically rely on fixed visual feature encodings extracted from vision foundation models, which are projected into the text space and passed to the language model along with instructions (Driess et al., 2023b; Liu et al., 2024a; b). However, the static nature of these encodings, formed without considering the instruction, limits the model’s ability to adapt dynamically to specific tasks or contexts. This disconnect between visual and textual components reduces flexibility, making the model less responsive to task-specific nuances. To address this, BLIP-2 (Li et al., 2023b) introduced a cross-attention-based module (called Q-former) to integrate the visual and instruction encodings, a design later adopted by others (Li et al., 2023b; Huang et al., 2023; Ye et al., 2023b). The current state-of-art model, mPLUG-Owl2 (Ye et al., 2024), introduces a modality adaptation module that employs attention modules to embed the two modalities into a shared semantic space and thus enhances cross-modality collaboration. mPLUG-Owl2’s performance improvement comes with several limitations compared to baseline models. First, mPLUG-Owl2 leverages LLaMA’s text embeddings and CLIP’s vision encoder to generate the instruction and visual encodings, respectively. Therefore, the encodings are generated using two distinct models with no initial multi-modality alignment. Second, mPLUG-Owl2’s modality-adaptive module introduces roughly 1B more parameters, 3×3$\times$3 × more than its vision encoder. The modularity adaptation module is then trained from scratch, requiring 348 million image-text pairs for pertaining, 300×300$\times$300 × more than the baseline. Third, the vision encoder requires training during both the pretraining and instruction-tuning stages, which increases the overall training cost and makes the vision encoder more susceptible to loss of generality. Finally, except for a few benchmarks, the model offers only marginal improvements, and in some cases, performs worse than the baseline.
@@
  • EMMA Outperforms Larger Models: Compared to mPLUG-Owl2, which has 50×50$\times$50 × larger modality adaptation module and is trained on 300×300$\times$300 × more data, EMMA outperforms on 7 of 8 benchmarks. Additionally, compared with BRAVE, which has 24×24$\times$24 × larger vision encoder and is trained on 100×100$\times$100 × more data, EMMA outperforms on all benchmarks.
@@ 2 Related Work
In recent years, there has been significant progress in the development of multi-modal large language models (LLMs) that integrate vision and language to handle tasks requiring both modalities (Zhang et al., 2024; 2023; Wu et al., 2023; Sun et al., 2024; Alayrac et al., 2022; Lai et al., 2023; Li et al., 2023a; g; d; Lin et al., 2024; Liu et al., 2023b; c; Tian et al., 2024; Wang et al., 2024b; c; Chen et al., 2023a). By combining the language understanding of LLMs with the perceptual abilities of vision foundation models, multi-modal LLMs are able to tackle a wide array of tasks that require cross-modal alignment and understanding. We can divide these models into two general categories based on the way that the two visual and textual modalities are integrated. In the first category, including LLaVA (Liu et al., 2024b; a), PaLM-E (Driess et al., 2023a), Shikra (Chen et al., 2023b), etc., the vision encodings are projected to the text space with a few linear layers, and then concatenated with the instruction tokens and passed to the LLMs. In the second category, a more complex module is used for cross-modality adaptation, where both visual and textual encodings are processed through the adaptation module. First introduced with Flamingo (Alayrac et al., 2022) and later adopted by BLIP-2 (Li et al., 2023c), InstructBLIP (Dai et al., 2024), Qwen-VL (Bai et al., 2023), mPLUG-Owl (Ye et al., 2023a), and MiniGPT-4 (Zhu et al., 2024), the use of a Q-former (a cross-attention-based module) has become a prominent technique.
@@
InstructBLIP 1.3B ×$\times\text{3.3}$ 200M×$\times\text{10}$ 7B 7.91B
BLIP-2/InstructBLIP 1.3B×$\times\text{3.3}$ 200M×$\times\text{10}$ 13B 14.2B
Qwen-VL-Chat 1.9B×$\times\text{6.3}$ 80M×$\times\text{4}$ 7.7B 9.6B
BRAVE 7.2B×$\times\textbf{24}$ 100M×$\times\text{5}$ 2.6B 10B
mPLUG-Owl2 0.3B×$\times\text{1}$ 1000M×$\times\textbf{50}$ 7B 8.2B
EMMA 0.3B 22M 7B 7.3B
Table 1: Parameter sizes of EMMA compared with leading MLLMs across three components: Vision Encoder, Modality Adapter, and LLM. EMMA uses the same-sized Vision Encoder and LLM as mPLUG-Owl2, but its Modality Adapter, our main contribution, is 50$\times$ smaller.
@@
Since the emergence of multi-modal LLMs, aligning visual and textual modalities has remained a significant challenge in achieving robust and seamless integration. Previous works (Alayrac et al., 2022; Li et al., 2023c; Ye et al., 2023a; c; Kar et al., 2024) have predominantly focused on utilizing Q-formers and similar cross-attention modules as modality adaptation components to integrate visual and textual embeddings. Recent advancements in modality adaptation include mPLUG-Owl2 (Ye et al., 2023c) and BRAVE (Kar et al., 2024). mPLUG-Owl2 (Ye et al., 2023c) introduces 1B modality adaptation module that employs distinct parameters to project multiple modalities into a unified semantic space for enhanced modality adaptation. On the other hand, BRAVE (Kar et al., 2024) leverages a concatenation of various visual encodings, summing up to 7B, directly feeding into the Q-former. These modality adaptation modules rely on intricate architectures, introducing millions to billions of additional parameters to the model, which significantly increases the computational costs. Recently, proposed a Question-Aware Vision Transformer approach that integrates question-awareness directly into the vision encoder. While this design enhances task-specific performance, it compromises the generality of the visual representations. This added complexity not only demands vast amounts of training data but also imposes considerable overhead during inference. Furthermore, the complication of these systems makes it difficult to discern the primary drivers behind performance improvements — whether they stem from the model complexity, the early fusion of vision and text encodings, or the sheer volume of additional training data. This leads us to the central focus of our work: (1) addressing the inefficiencies in current multi-modal models by proposing a more streamlined approach. Our goal is to achieve an efficient early fusion of visual and textual encodings without significantly increasing the number of parameters or computational costs. (2) Conducting a thorough analysis to determine the key drivers behind performance improvements and the dynamics of the modality adaptation module to generate the multi-modal encodings. This investigation will help isolate the most effective factors and guide future advancements in multi-modal LLMs.
@@
The evaluation of multi-modal LLMs relies on a mix of traditional academic benchmarks and newer ones tailored to instruction-following MLLMs. Established benchmarks like VQA-v2 (Goyal et al., 2017) and GQA Hudson & Manning (2019) gauge a model’s ability to interpret visuals through open-ended, short-answer questions. ScienceQA (Lu et al., 2022b) tests zero-shot generalization in scientific question answering, while VizWiz (Gurari et al., 2018) offers real-world images captured by visually impaired users, challenging models with poor framing, blur, and other quality issues typical of non-professional photos. Additionally, newer benchmarks target instruction-following MLLMs. MathVista (Lu et al., 2023) introduces diverse challenges from mathematical and visual tasks. MMMU (Yue et al., 2024) evaluates multi-modal models on a broad range of college-level tasks requiring deep subject knowledge and reasoning. To assess multi-image reasoning, MUIRBENCH (Wang et al., 2024a) provides a comprehensive benchmark with 12 diverse tasks to evaluate the multi-image understanding abilities of MLLMs.
For general robustness, MMBench (Liu et al., 2023a) offers a multiple-choice visual question answering benchmark in both English and Chinese, suggesting shuffling the choices to test the model’s robustness to the order of options. MMVP (Tong et al., 2024) evaluates robustness by identifying similar images with minute differences and manually pinpointing the visual details the CLIP vision encoder overlooks, which leads to incorrect responses from MLLMs. For hallucination, POPE (Li et al., 2023f) examines the extent of hallucinations across three COCO subsets, and AMBER (Wang et al., 2023b), a multi-dimensional benchmark, evaluates generative and discriminative tasks. Additionally, three benchmarks were utilized to assess the proposed method’s robustness against hallucination: FOIL (Shekhar et al., 2017), MMRel (Nie et al., 2024), and R-Bench (Wu et al., 2024).
@@ 3 Efficient Visual Alignment in Multi-Modal LLMs
Figure 1: This figure presents the EMMA’s Architecture where Modality Adaptation is introduced to the standard MLLM architecture. The Modality Adaptation module consists of the Text Encoder, Instruction Projection, and Visual alignment modules. EMMA enhances the multi-modality alignment by incorporating the instruction encodings generated by the CLIP’s Text Encoder, which are then projected to the joint space by the Instruction Projection module. The concatenating visual and textual encodings are then passed through the Visual Alignment module, resulting in multi-modal, instruction-aware representations.
@@
Figure 2: To evaluate the contribution of visual and textual tokens to the Aligned Representations, we analyze the weight matrix of the Visual Alignment module, which consists of a linear layer. Left: Fig. 2(a) presents the $\ell_{1}$ norms of the weights for each token, revealing that (1) visual tokens have a stronger influence on the aligned representations and (2) the impact of textual tokens varies, prompting further investigation in Fig. 2(b) to identify which specific textual tokens contribute more significantly to the final representations. Right: The bar plot reveals that the early tokens are assigned higher weights in the Visual Alignment module, placing greater emphasis on them.
@@ 3.1 EMMA: Efficient Multi-Modal Adaptation
Recent progress in multi-modal models has been largely driven by the robust reasoning capabilities of large language models. Therefore, a persistent challenge is to effectively align these two modalities to ensure seamless fusion and task-specific adaptability. Current approaches often rely on complex cross-modality modules, which introduce a significant number of parameters and thus require large amounts of training data, as shown in Table 1. We hypothesize that the need for a complex modality adaptation module arises from the fact that visual and textual encodings are produced by two independently trained components that are themselves unaligned. This is precisely the case for mPLUG-Owl2, which uses CLIP as its vision encoder and LLaMA’s text embeddings for its text encoder. As a result, the multi-modality module must, in addition to incorporating text information into the visual embedding, also align the two embeddings.
@@
We refer to our proposed architecture as EMMAEfficient Multi-Modal Adaptation. Figure 1 illustrates EMMA’s architecture. The left side highlights the high-level structure, where the standard modules of a multi-modal LLM are shaded in gray, while EMMA’s newly introduced Modality Adaptation module is shown in green. On the right, the details of the Modality Adaptation module are depicted. More specifically, let $v(\cdot)$ and $t(\cdot)$ represent the vision encoder and text encoder of CLIP, respectively. The visual encodings are defined as $\mathbf{v}=v(\mathbf{x})\in\mathbb{R}^{n\times d}$, where $\mathbf{x}$ is the input image, and the text encodings as $\mathbf{t}=t(\mathbf{i})\in\mathbb{R}^{m\times d^{\prime}}$, where $\mathbf{i}$ is the instruction. Later, the instruction encodings are then processed through the Instruction Projection module $p:\mathbb{R}^{d^{\prime}}\rightarrow\mathbb{R}^{d}$, which maps the instruction tokens to the same dimensional space as the visual tokens. Once the visual and textual representations are generated, we introduce early fusion via a lightweight module called the Visual Alignment module. This component, consisting of a simple linear layer, combines visual and textual tokens to create the model’s multi-modal encoding. The Visual Alignment module, which consists of a linear layer, forms the core of our proposed method. Let $f:\mathbb{R}^{(n+m)\times d}\rightarrow\mathbb{R}^{n\times d}$ represent the Visual Alignment module, where $n$ and $m$ denote the number of visual and textual tokens, respectively. The Visual Alignment module takes the concatenation of the visual and textual tokens as input to generate $n$ refined visual tokens $\tilde{\mathbf{v}}$. The complete pipeline can be expressed as follows:
$\tilde{\mathbf{v}}=f(v(\mathbf{x}),\;p(t(\mathbf{i}))$ (1)
Note that the dimensions of this alignment layer are designed to maintain the same number of visual tokens as the baseline model, ensuring consistency in the number of visual tokens passed to the language model. The Visual Alignment module plays a critical role in ensuring effective alignment between visual and instructions encodings, highlighting the most relevant tokens from the visual encodings in response to the instruction, and thereby delivering more precise visual information to the language model. Its lightweight design facilitates (1) easier interpretability and analysis, and more importantly, (2) its simplicity mitigates overfitting when training on small datasets and (3) reduces the training and inference time while outperforming the state-of-the-art models with 10×10$\times$10 × fewer parameters. Moreover, by leveraging the inherent alignment of CLIP’s vision and text encoders, both encoders remain frozen during training. This prevents overfitting to the training data and preserves the generalization capabilities of the original encoders, enabling EMMA to deliver robust performance across diverse tasks.
Figure 3: Left: Figure 3(a) illustrates the mutual information between visual representations and response encodings. EMMA’s visual representations exhibit 1.5$\times$ higher mutual information with the response encodings compared to LLaVA’s, highlighting its superior alignment with the instruction. Right: Figure 3(b) isolates the impact of EMMA’s modality adaptation module by comparing it to the baseline model, LLaVA-1.5, while keeping all other parameters constant except for the architectural innovation.